OpenStudy (anonymous):

solve the equation 7x+10=-18

OpenStudy (anonymous):

1-Subtract 10 from both sides . 2-divide by 7 .

OpenStudy (anonymous):

x= -4

OpenStudy (anonymous):

thnak you guys

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

7x=-18-10 7x=-28 x=-28/7 x=-4